OpenStudy (anonymous):

Solve this Separable Equation: x(dy/dx)=1/y^3 with explanations for each step please

OpenStudy (anonymous):

x(dy/dx) = 1/y^3 - for example x * dy = dx / y^3 - communitive property y^3 * x dy = dx - these aren't right y^3 dy = dx/x - but explain each one y^4/4 = ln(x) + C - my teacher is really picky y^4 = 4*ln(x) + 4C - and i dont remember all the terms and lingos y^4 = 4*[ln(x) + C] - integration by parts??? please help

OpenStudy (mathmale):

Your result is fine. I'd do it slightly differently, writing (1/4)y^4=ln|x|+ln C = (1/4)y^4=ln |Cx| (which I think is simpler and less trouble-prone). Solving for y^4, multiply both sides of the equation by 4: y^4=4 ln |Cx| = ln (Cx)^4 Then, solving for y, \[y=(\ln (Cx))^{\frac{ 1 }{ 4 }}\]

OpenStudy (mathmale):

I see you wanted explanations of some of the steps involved. Let me know which ones that applies to.

OpenStudy (anonymous):

all of them if possible, my prof doesn't mind if they are simple steps like dividing by 4, but integrating and such

OpenStudy (anonymous):

quick Q how did the 4 go away in front of ln?

OpenStudy (mathmale):

\[dx(x \frac{ dy }{ dx }=\frac{ 1 }{ y ^{3} })\rightarrow xdy = \frac{ dx }{ y ^{3} } commutative property\] Rearrange so that all x components are together and all y comps are together; I belileve this is also due to the commutative property: multiply the entire equation through by (y^3)/x

OpenStudy (mathmale):

Please type back to me what you now have...what does our equation look like?

OpenStudy (anonymous):

y=(ln(Cx))14

OpenStudy (anonymous):

y=(ln(Cx))^1/4

OpenStudy (mathmale):

You're a couple of steps too far ahead. What I was asking you to do is the following:

OpenStudy (anonymous):

x(dy/dx) = 1/y^3 x * dy = dx / y^3 y^3 * x dy = dx y^3 dy = dx/x y^4/4 = ln(x) + C y^4 = 4*ln(x) + 4C y^4 = 4*[ln(x) + C] y = sqrt(2) * [ln(x) + C]^(1/4) or y = -sqrt(2) * [ln(x) + C]^(1/4) thats what i got

OpenStudy (mathmale):

\[(\frac{ y^3 }{ x })(x \frac{ dy }{ dx }=\frac{ 1 }{ y^3 }).\]Please do that and type back your result.

OpenStudy (mathmale):

Sorry, I meant to ask you to multiply by \[\frac{ y^3dx }{ x }.\]

OpenStudy (anonymous):

im lost im sorry

OpenStudy (mathmale):

\[\frac{ y^3dx }{ x }(\frac{ dy }{ dx }=\frac{ 1 }{ y^3 })\rightarrow y^3dy=\frac{ dx }{ x }.\]

OpenStudy (mathmale):

Make sure you understand this before we move on. Which algebraic property did we use here?

OpenStudy (anonymous):

\[x(y ^{3}dy/x dx)=y^{3}dx/xy^{2}\]

OpenStudy (anonymous):

commutative i believe

OpenStudy (mathmale):

that's right. Compare your result to mine. You'll see that in mine, all x are grouped together in one term and all y together in another term. This is critically important. You are approaching the sol'n of this problem using the "separation of variables" method.

OpenStudy (mathmale):

IF:\[y^3dy=\frac{ dx }{ x },\] our next step is to integrate both sides so as to obtain an expression for \[\frac{ y^4 }{ 4 }.\]

OpenStudy (mathmale):

Remember, we started out with an equation that has derivatives in it and want to end up with an equation in x and y that does NOT have derivatives in it. This is the basic thesis of Differential Equations.

OpenStudy (mathmale):

OK with that? If not, ask questions now, please.

OpenStudy (anonymous):

I completely get it

OpenStudy (anonymous):

so y^4/4=ln|x|+c

OpenStudy (mathmale):

So now we do the actual integration:\[\int\limits_{-}^{-}y^3dy=\int\limits_{-}^{-}\frac{ dx }{ x }\]

OpenStudy (anonymous):

why dont you put dx at the end of it?

OpenStudy (mathmale):

Use the power rule of integration to integrate the left side; on the right side we have a special case in which the appropriate result is ln |x| + C or len |x| + ln D or ln D|x}.

OpenStudy (anonymous):

reciprocal rule for the right

OpenStudy (mathmale):

Actually, I did: \[\int\limits_{-}^{-}\frac{ 1 }{ x }dx \] is exactly the same as what I wrote before. that's not "reciprocal rule," but rather a special case of the power rule for integration.

OpenStudy (mathmale):

\[\int\limits_{-}^{-}\frac{ dx }{ x }= \ln |x| + c\]

OpenStudy (anonymous):

on my paper i wrote integral sign dx/ x dx to close

OpenStudy (mathmale):

Tell me when you're ready to move forward.

OpenStudy (anonymous):

i am ready

OpenStudy (mathmale):

Integrating both sides, we end up with:\[\frac{ y^4 }{ 4 }=\ln |x| + C, or \frac{ y^4 }{ 4 }=\ln |x| + \ln D= \ln D|x|\]

OpenStudy (mathmale):

Each time I type something new, please respond with "I am ready to continue" or ask a pertinent question.

OpenStudy (anonymous):

gotcha

OpenStudy (mathmale):

Then:\[\frac{ y^4 }{ 4 }=\ln D|x|\]

OpenStudy (mathmale):

WHICH can be multiplied by 4 (distributive property) to obtain:

OpenStudy (anonymous):

"I am ready to continue"

OpenStudy (mathmale):

\[y^4=4\ln D|x|=\ln (D|x|)^4\]

OpenStudy (mathmale):

What do you think we need to do next?

OpenStudy (anonymous):

fourth root everything

OpenStudy (mathmale):

Because of the "ln" in there, that would not be appropriate. We could leave the last result as is, or we could eliminate the "ln". Do you know how to evaluate \[e^\ln x\]??

OpenStudy (mathmale):

(ignore that " \ " character, please)

OpenStudy (anonymous):

isn't that just x

OpenStudy (mathmale):

Right. very good!

OpenStudy (mathmale):

Therefore, if:\[y^4=\ln(D|x|)^4,\]

OpenStudy (mathmale):

use each side of this equation as the exponent of e (raise e to each of these 2 "powers"):

OpenStudy (mathmale):

\[\exp y^4=\exp \ln (D|x|)^4\]

OpenStudy (mathmale):

then, on the left side, you'd have e^(y^4), and on the right side just (D|x|)^4. End of solution.

OpenStudy (mathmale):

Unless, of course, you want to solve for y; in that case you'd just take the 4th root of both sides of the equation I've just given you. Satisfied with the explanations of the steps or have you further qeustions?

OpenStudy (anonymous):

so e^y^4=D|x|^4

OpenStudy (mathmale):

"exp" signifies the use of e as the base: exp 5 translates to d^5.

OpenStudy (anonymous):

so lets solve for y

OpenStudy (mathmale):

All y ou have to do then is to take the fourth root of each side of my last equation;\[(y^4)^{1/4}=((D|x|)^4)^{1/4}\]

OpenStudy (anonymous):

this answer is way different than the first one you wrote

OpenStudy (mathmale):

yES, i realize that. Can YOU fix it?

OpenStudy (anonymous):

ouch

OpenStudy (anonymous):

no

OpenStudy (mathmale):

Ouch is right. But great practice for you if you can identify my error and fix it. :)

OpenStudy (anonymous):

i;m going to cry, this is my first semester of DE

OpenStudy (anonymous):

was this all just a game to you

OpenStudy (mathmale):

Aw, come on! Look:

OpenStudy (mathmale):

\[\exp y^4=\exp \ln(D|x|)^4 becomes \exp y^4=(D|x|)^4.\]

OpenStudy (mathmale):

Personally, I'd leave the answer like that.

OpenStudy (mathmale):

One more time: any final question or questions?

OpenStudy (anonymous):

Then, solving for y, y=(ln(Cx))14

OpenStudy (anonymous):

that is what you wrote as the answer way up above

OpenStudy (mathmale):

Certainly not just a game to me, but I'm a retired math professor and have been through these steps countless times.

OpenStudy (anonymous):

you wrote "Then, solving for y, y=(ln(Cx))^1/4"

OpenStudy (mathmale):

I'll try to answer specific questions and am sorry I ended up with an answer different from my previous one, but do hope you will go through all this on your own and make up your own mind which result is the correct one. Eventually you'll have to do all of the work on problems such as these on your own, so the more experience you get with them, the better.

OpenStudy (mathmale):

Jamal, it does seem that you have a point. Earlier I wrote that the final answer was something like \[y=fourth \root of \ln (D|x|)^4\], and after having done the problem again, that's the result I got. Apologies. But at this point you 'll have to decide what YOU want to do next. Start from scratch? or move on to another problem? or?

OpenStudy (anonymous):

im going to cry

OpenStudy (mathmale):

Sorry. Wish I could give you a big comfort hug. But I can't, not from California, anyway. So: Please choose one of the alternatives I've given you.

OpenStudy (anonymous):

this is the original equation x(dy/dx)=1/y^3

OpenStudy (mathmale):

Yes, I'm doing the whole problem over again on paper. But before that, could you possibly look through what we've done with a critical eye and decide which part, i f any, makes sense to you? It's not just the answer that's important, but the process of arriving at it.

OpenStudy (anonymous):